The points (-1,3) and (0,2) are on a curve, and at any point (x,y) on the curve ๐‘‘ 2๐‘ฆ ๐‘‘๐‘ฅ 2 = 2 โˆ’ 4๐‘ฅ. Find an equation of the curve.ย 


1
Expert's answer
2021-03-01T07:11:06-0500
"\dfrac{d^2y}{dx^2} = 2-4x\\
d^2y =(2-4x)dx^2\\
d(dy) = [(2-4x)dx]dx\\
\int d(dy) = \int[(2-4x)dx]dx\\
dy = (2x - 2x^2 +c_1)dx\\
\int dy = \int(2x - 2x^2 +c_1)dx\\"

The general solution of the curve is given as:


"{y =x^2-\frac{2x^3}{3}+c_1x + c_2}"

at point (-1,3):


"3 = (-1)^2 - \frac{2(-1)^3}{3} +c_1(-1) + c_2\\
3 = 1 + \frac{2}{3} - c_1+c_2\\
3-1-\frac{2}{3} = -c_1+c_2
\implies \frac{4}{3} = -c_1+c_2"

At point (0,2):


"2 = (0)^2 - \frac{2(0)^3}{3} +c_1(0) + c_2\\
\implies c_2=2"

Substitute for "c_2" in the previous equation:


"\frac{4}{3} = -c_1+2\\
c_1 = 2 - \frac{4}{3}\\
c_1 = \frac{2}{3}"

"\therefore" the equation of the curve is:


"y =x^2-\frac{2x^3}{3}+\frac{2x}{3} + 2 \implies y =2+\frac{2x}{3}+x^2-\frac{2x^3}{3}"